Determine the permittivity of a dielectric within a capacitor stores 60 Mf worth or charge has an area of 0.5 m² for each plate and 2.0 X 10^-⁴ m apart.


1
Expert's answer
2021-05-03T10:31:44-0400

Answer

Capacitance is given by

"C=\frac{\epsilon _0KA}{d}"

K is permittivity of dielectric medium.

"K=\frac{Cd}{ A \epsilon _0 }\\=\frac{60\times10^{-6}\times 2.0 \times10^-⁴ }{ 0.5\times8.85\times10^{-12}}\\=2711.86"